OpenStudy (anonymous):

HELP! sin x cos x = 1/2 all solutions [0, 2pi]

OpenStudy (anonymous):

sin x = 1/2 cos x = 1/2 x = pi/6, 5pi/6 (sin) x = pi/3 , 5pi/3 (cos)

OpenStudy (anonymous):

My answer sheet says pi/4, 5pi/4 any formula i can put it in to get that?

OpenStudy (anonymous):

oh okay okay sorry

OpenStudy (anonymous):

so when you have a unit circle, you can see that at pi/4, its going to be (rad(2)/2 , rad(2)/2)

OpenStudy (anonymous):

got that part? and when you multiply them, you get 1/2

OpenStudy (anonymous):

but, there is another answer. When you multiply the negatives together, you get positive.

OpenStudy (anonymous):

so since at quadrant 3, 5pi/4 is (-rad(2)/2, -rad(2)/2), you still are going to get 1/2.

OpenStudy (anonymous):

so that's why the answer is pi/4 and 5pi/4.

OpenStudy (anonymous):

thank you! much appreciated

OpenStudy (anonymous):

no problem :)